# Improved Belief Propagation Decoding on Surface Codes with High Accuracy and Low Latency

### v1

@@ Abstract
Quantum error correction is crucial for universal quantum computing, requiring highly accurate and low-latency decoding algorithms. Belief Propagation (BP) is notable for its linear time complexity and general applicability to quantum LDPC codes. However, BP performs poorly on highly degenerate codes without Order Statistic Decoding (OSD) post-processing, which significantly increases time complexity. We focus on improving BP’s performance on surface codes. We first propose Momentum-BP and AdaGrad-BP, inspired by machine learning optimization techniques, to reduce the oscillation of message updating and break the symmetric trapping sets. We further propose EWAInit-BP, which adaptively updates initial probabilities and exhibits aggressive exploration capabilities. EWAInit-BP achieves the highest accuracy among BP improvements without OSD post-processing on planar surface code, toric code, and XZZX surface code, providing a 1~3 orders of magnitude improvement compared to traditional BP, and demonstrating the error correction capability even under parallel scheduling. Its theoretical $O(1)$ time complexity and high accuracy make it a promising candidate for high-precision real-time decoders.
1 Introduction
@@
Information in quantum computing devices is highly susceptible to noises and interference, leading to information loss and computational errors [1]. To address this issue, quantum error correction (QEC) technology introduces redundant qubits and uses quantum error-correcting codes to protect the integrity of quantum information. In fact, QEC is an essential step towards achieving fault-tolerant quantum computation [2].
Most existing quantum error correction codes are based on stabilizer framework [3]. Surface codes are a special class of stabilizer codes where qubits are placed on a 2D lattice [4]. These codes, especially the planar surface code have received significant attention over the past two decades [5, 6, 7] due to the local connections between stabilizers and qubits and high circuit-level threshold [8]. The planar surface code is defined on the lattice with open boundaries, which can reduce engineering requirements [9]. The toric code, on the other hand, is encoded in a torus-like structure with periodic boundaries, featuring complete boundary stabilizers and thus slightly stronger error-correcting capabilities [10]. The XZZX surface code is an improvement over the planar surface code, offering strong performance under biased noise [11]. All the three codes have the code distance scaling with the code length as $O(\sqrt{N})$ but have asymptotically vanishing code rates.
There are several decoding algorithms suitable for surface codes. The Minimum-Weight Perfect Matching (MWPM) algorithm [12] achieves a code capacity threshold of approximately 15.5% for the planar surface code under depolarizing noise, but its time complexity is $O(N^{3})$. The Tensor Network Contraction method [13] achieves a very high threshold of 18.3%, but its time complexity is $O(N\gamma^{3})$, where $\gamma$ controls the approximation accuracy. The Renormalization Group method [14] achieves a threshold of 15.2%, with the time complexity of $O(N\log{\sqrt{N}})$. The Union-Find method [15] has near-linear complexity $O(N\alpha)$, but it achieves a slightly lower threshold of 14.8%. However, the aforementioned methods can hardly simultaneously meet the requirements for high accuracy and low latency in quantum error correction decoding [16, 17].
@@
Belief Propagation (BP) algorithm [18] are applicable to any sparse quantum code and can achieve a linear complexity of $O(j\tau N)$, where $j$ is the average weight of the rows/columns of the check matrix, and $\tau$ is the number of iterations. Additionally, parallel scheduling of BP can be efficiently implemented in hardware, theoretically achieving constant-time complexity [19]. However, traditional BP exhibit low accuracy for many quantum error correcting codes and even cannot achieve the code capacity threshold for surface codes. BP can be implemented in both quaternary [20, 21, 22, 23, 24, 25, 26] and binary domain [27, 28, 29, 30]. There have been some improvements that enable BP to demonstrate error-correcting capabilities on surface codes (i.e., logical error rates decrease as code length increases). In the quaternary domain, MBP [25] achieves a threshold of 14.5%~16% on the planar surface code by modifying the step size in the a posteriori probability updating. However, MBP can only implemented in serial scheduling, and the outer loop for searching step sizes increases its time complexity. In the binary domain, GBP [30] achieves a 14% threshold by combining generalized belief propagation with multiple rounds of decoding. This algorithm can be implemented in parallel, but the number of outer loops increases with the code length, adding to its time complexity. BP-OTS [31] periodically detects and breaks trapping sets using manually set parameters, allowing parallel implementation while maintaining the original complexity. It outperforms MBP for short codes but does not achieve a significant threshold.
BP can also be combined with post-processing methods, such as BP-MWPM [32, 33, 34], BP-SI (stabilizer inactivation) [35]. Although these algorithms can achieve high thresholds for surface codes, they all have high time complexities, ranging from $O(N^{2})$ to $O(N^{3})$. For example, BP-OSD [36, 37, 38] is renowned for its high thresholds on most sparse quantum codes, particularly achieving 17.68% on planar surface codes [39]. However, even using the lowest 0-order OSD, the time overhead in hardware is unacceptable for real-time decoding [19]. Enhancing BP performance without relying on post-processing algorithms remains a significant challenge.
@@
There are several explanations for why surface codes are challenging for BP decoding. From the perspective of degeneracy [40, 25], stabilizers of surface codes have much lower weights compared to the code distance, leading to multiple low-weight errors corresponding to the same syndrome, thus complicating BP’s error convergence. From the perspective of short cycles [18], a number of 4-cycles in the quaternary domain and 8-cycles in the binary domain in surface codes can cause BP’s inaccurate updating. A more intuitive explanation involves trapping sets [41, 31], where surface codes have many symmetric trapping sets causing BP messages to oscillate periodically.
@@
This paper employs three strategies for improving BP: optimizing message updates, optimizing prior probabilities, and utilizing iterative feedback information. First, we leverage the similarity between BP posterior updating and gradient descent of the energy function, drawing inspiration from optimization methods in machine learning to enhance the message update rules. In addition to reducing the logical error rate of decoding, experiments on trapping sets in surface codes show that our proposed method, Momentum-BP and AdaGrad-BP, can smooth the message update direction and reduce oscillation amplitude.
@@
Furthermore, recognizing the importance of prior probabilities from the probabilistic derivation of BP, we extend the idea of optimizing message updates to an adaptive approach, EWAInit-BP, to dynamically update the a priori probabilities involved in each iteration. Experiments on short toric codes, planar surface codes, and XZZX surface codes (under bias noise) show that EWAInit-BP outperforms existing BP improvements without post-processing while demonstrating error correction capabilities on both parallel scheduling and serial scheduling. Notably, the time complexity of all proposed improvements remains $O(j\tau n)$. Thus, EWAInit-BP can be implemented with theoretical constant complexity in parallel while maintaining decoding accuracy.
@@
The structure of this paper is as follows. Section 2 introduces the fundamentals of surface codes and BP decoding, as well as the perspective of BP as a gradient optimization, and summarizes the challenges BP faces in decoding surface codes. Section 3 first optimizes BP’s message update rules, and then applies these ideas to the adaptive initialization. The results of the simulation are presented in Chapter 4. Finally, Chapter 5 summarizes our work and discusses the existing challenges.
2 Preliminaries
2.1 Surface codes
@@
A stabilizer code is defined by an abelian subgroup $\mathcal{S}$ of the Pauli group $\mathcal{P}_{n}$ excluding $-I$, composed of operators $S\in\mathcal{P}_{n}$ such that $S|\psi\rangle=|\psi\rangle$ for all codewords $|\psi\rangle$ in the code space. The code space is the simultaneous $+1$ eigenspace of all elements of $\mathcal{S}$. The stabilizers must commute with each other, which means $[S_{i},S_{j}]=0$ for all $S_{i},S_{j}\in\mathcal{S}$. The dimension of the code space is determined by the number of independent generators of $\mathcal{S}$. For a stabilizer code with $N$ physical qubits and $K$ logical qubits, $\mathcal{S}$ has $N-K$ independent generators.
@@
The stabilizers are used to detect errors by measuring their eigenvalues. The set of measurement outcomes, known as the error syndrome, indicates the presence and possibly the location of errors. Logical operators are Pauli operators that commute with all elements of $\mathcal{S}$ but are not themselves in $\mathcal{S}$, thus in $\mathcal{N(S)\backslash S}$ where $\mathcal{N(S)}=\{E:ES=SE,\ \forall S\in\mathcal{S}\}$. They act on the encoded logical qubits without disturbing the code space, thus cause an undetectable error. The code distance $D$ is defined as the minimum weight of logical operators. Therefore, a quantum stabilizer code can be represented by $[[N,K,D]]$.
@@
Surface codes are a class of quantum Low-Density Parity-Check (LDPC) codes [42] where qubits are placed on a 2-dimensional lattice, relying on the topology of the surface to protect quantum information [4]. Logical qubits are encoded in global degrees of freedom, and larger lattices provide larger code distances. There are some common properties between different types of surface codes, such as the code distance scales with the code length as $O(\sqrt{N})$, while the code rate is asymptotically zero.
@@
The $[[2L^{2},2,L]]$ toric code has a 2-dimensional torus topology, and the closed boundary makes it translationally invariant to the error syndrome, which is beneficial to decoding algorithms (such as BP and neural networks). The Euler characteristic of the torus is 0, but the X and Z stabilizers of the toric code each have a redundancy, resulting in a total of two encoded logical qubits.
@@
The $[[2L^{2}-2L+1,1,L]]$ planar surface code is defined on a surface with open boundaries, which makes the stabilizers near the boundaries having lower weights. This reduction in stabilizer weight results in a slight decrease in the error correction capability compared to the toric code, however, simplifies its implementation in physical quantum systems. The lattice structure of the planar surface codes are illustrated in Fig. 1.
@@
The $[[L^{2},2,L]]$ XZZX surface code is a non-CSS code designed to handle biased noise more effectively. The stabilizers of the family of code have the form of XZZX. This alternating pattern of Pauli-X and Pauli-Z operators allows the code to reach the theoretical maximum threshold of 50% for certain types of noise.
@@
Figure 1: The lattice of the $[[2L^{2}-2L+1,1,L]]$ planar surface code for L=3. Circles represent data qubits, yellow squares represent X-type stabilizers, green squares represent Z-type stabilizers, and arrows indicate the coupling between stabilizers and qubits.
2.2 Belief propagation decoding
2.2.1 BP in the quaternary logarithmic domain
@@
BP algorithm passes messages on a Tanner graph, where stabilizers and qubits are represented as check nodes and variable nodes, respectively. In the quaternary representation which this paper mainly focuses on, each variable node corresponds to a probability vector of all Pauli errors on one qubit.
The quaternary logarithmic domain belief propagation (abbreviated as LLR-BP4) receives a priori probability vectors describing each type of Pauli error on each qubit $Q^{(0)}=\{p_{v_{j}}^{X},p_{v_{j}}^{Y},p_{v_{j}}^{Z}\}_{j=1}^{N}$, and initializes the variable-to-check messages as [25]
@@
In the horizontal updates, each check node calculates the feedback based on the messages received from the variable nodes but not relying on the target variable node, sending a check-to-variable message
@@
In the vertical updates, variable nodes aggregate the messages from check nodes and, after a similar extrinsic calculation, send variable-to-check messages
@@
Then, for each variable node, if all log-likelihood ratios (LLR) for Pauli errors are greater than 0, then the hard decision result is I; otherwise, it is the Pauli error with the smallest LLR value. The hard decisions of all qubits combine to form the error estimate $\hat{E}$. If $\hat{E}$ corrects the error syndrome, the algorithm converges; otherwise, it returns to the horizontal update for a new iteration. If the maximum number of iterations is reached, decoding is considered to have failed.
BP can be performed using either parallel or serial scheduling [44]. In parallel scheduling, all variable nodes and check nodes update their messages simultaneously at each iteration, allowing for hardware parallelization, potentially achieving constant complexity. On the other hand, the serial scheduling updates the messages of variable nodes and check nodes sequentially. This method leads to faster convergence and slightly higher accuracy, though it has a more complex hardware implementation.
2.2.2 BP as gradient optimization
@@
From this gradient optimization perspective, MBP [25] adjusts the "learning rate" of message updates to converge more quickly to the equivalent energy minimum, which is common in degenerate quantum codes.
2.2.3 Hardness of BP decoding on surface codes
BP without post-processing algorithms perform poorly on surface codes due to factors from multiple perspectives.
@@ 2.2.3 Hardness of BP decoding on surface codes
Short cycles cause the joint probability calculations in BP to be dependent, which violates the independence assumption of updating variable nodes Eq. (7) and check nodes Eq. (5) [18]. Surface codes exhibit 8-cycles in the binary representation and 4-cycles in the quaternary representation, exacerbating this issue.
@@
Optimizing BP from the first two perspectives involves modifying the underlying principles of the algorithm. This paper will focus on trapping sets and seek methods to break the symmetry in message passing.
@@
3 Belief Propagation Optimization
Surface codes violate the assumptions of BP’s objective function and probability independence, making BP decoding difficult to converge. These issues are manifested in symmetric trapping sets, where the local structure of each node is identical, and message updates are symmetric under specific error patterns. Breaking the former symmetry involves modifying the code structure, while in this section, we focus on breaking the latter, the symmetry of message updates. Notably, the algorithms proposed in this chapter preserve the same time complexity as traditional belief propagation.
3.1 Message updates optimization
3.1.1 Update direction smoothing

### v2

# Improved Belief Propagation Decoding algorithms for surface codes

@@ Abstract
Quantum error correction is crucial for universal fault-tolerant quantum computing. Highly accurate and low-time-complexity decoding algorithms play an indispensable role in making sure quantum error correction works. Among existing decoding algorithms, belief propagation (BP) is notable for its nearly linear time complexity and general applicability to stabilizer codes. However, BP’s decoding accuracy without post-processing is unsatisfactory in most situations. This article focuses on improving the decoding accuracy of BP over GF(4) for surface codes. We first propose Momentum-BP and AdaGrad-BP, inspired by machine learning optimization techniques, to reduce oscillation in message updating and break the symmetric trapping sets. We further propose EWAInit-BP, which adaptively updates initial probabilities and provides a 1 to 3 orders of magnitude improvement over traditional BP for planar surface code, toric code, and XZZX surface code without any post-processing method, showing high decoding accuracy even under parallel scheduling. The theoretical $O(1)$ time complexity under parallel scheduling and high accuracy of EWAInit-BP make it a promising candidate for high-precision real-time decoders. Meanwhile, the ideas of the Momentum-BP, AdaGrad-BP and EWAInit-BP provide promising approaches to improve the decoding accuracy of BP to get rid of its reliance on post-processing.
preprint: APS/123-QED
I Introduction
@@
Information in quantum computing devices is highly susceptible to noises, leading to information loss and computational errors [1]. To address this issue, quantum error correction (QEC) introduces redundant physical qubits and encodes the information in certain subspace of the state space of all the physical qubits to protect the information. So far at least, QEC is an essential step towards universal fault-tolerant quantum computation [2, 3, 4]. To make sure QEC work, the decoder of a QEC code should output the estimation of error fast and accurate according to the error syndromes.
Several decoding algorithms have achieved notable success, including the minimum-weight perfect matching (MWPM) [5] algorithm, union-find (UF) algorithm [6], maximum likelihood (ML) algorithm [7], matrix-product-states-based (MPS-based) algorithm [7], renormalization group algorithm [8], and neural-network-based (NN-based) algorithm [9, 10, 11, 12, 13, 14, 15, 16]. Among these, ML, MWPM, and MPS-based algorithms offer high accuracy but come with high complexity. In contrast, the UF algorithm provides near-linear complexity with slightly lower accuracy. NN-based decoders have to balance scalability, decoding accuracy, and time cost [13], and may struggle with generalizability across different codes. Consequently, existing decoding algorithms rarely meet the requirements for both high accuracy and low complexity in quantum error correction decoding [17, 18].
@@
Belief Propagation (BP) algorithm [19] are applicable to any stabilizer code and can achieve a complexity of $O(j\tau N)$ [20], where $j$ is the average weight of the rows/columns of the check matrix, and $\tau$ is the number of iterations. Additionally, parallel scheduling of BP can be efficiently implemented in hardware, theoretically achieving constant-time complexity [21]. However, traditional BP exhibit low decoding accuracy for many quantum error correcting codes and even cannot achieve the code capacity threshold for surface codes.
BP can be implemented over both GF(4) [22, 23, 24, 25, 26, 20, 27] and GF(2) [28, 29, 30, 31], where the Pauli operators are represented. There have been some efforts improving decoding accuracy of BP over GF(4) and GF(2). Over GF(4), MBP [20] achieves a threshold of 14.5% to 16% of the planar surface code by modifying the step size in the a posteriori probability updating. However, MBP is limited to serial scheduling, and the outer loop for searching step sizes increases its time complexity. Over GF(2), GBP [31] achieves a 14% threshold by combining generalized belief propagation with multiple rounds of decoding. This algorithm can be implemented in parallel, but the number of outer loops increases with the code length, adding to its time complexity. BP-OTS [32] periodically detects and breaks trapping sets using manually set parameters, allowing parallel implementation while maintaining the original complexity. It outperforms MBP for short codes but does not achieve a significant threshold.
Post-processing methods, such as OSD, SI (stabilizer inactivation) [33], MWPM [34, 35, 36] and UN[3] are often used to improve BP’s decoding accuracy. However, they also introduce extra time cost and all have high time complexities, ranging from $O(N^{2})$ to $O(N^{3})$. Improving BP’s decoding accuracy to get rid of its reliance on post-processing algorithms remains a significant challenge.
@@
In this article, we mainly focus on improving BP’s decoding accuracy for surface codes. However, we believe that the idea used in these improvements will also work for some other stabilizer codes. There are several explanations for why surface codes are challenging for BP decoding. From the perspective of degeneracy [37, 20], stabilizers of surface codes have much lower weights compared to the code distance, leading to multiple low-weight errors corresponding to the same syndrome, thus having negative influence on BP’s error convergence. From the perspective of short cycles [19], a number of 4-cycles over GF(4) and 8-cycles over GF(2) in surface codes can cause BP’s inaccurate updating. A more intuitive explanation involves trapping sets [38, 32], where surface codes have many symmetric trapping sets causing BP messages to oscillate periodically.
@@
This article employs three strategies to improve BP: optimizing message updates by optimization methods in machine learning, and optimizing a priori probabilities by utilizing a posteriori probabilities output by previous iterations. First, inspired by the similarity between BP message update and gradient descent of the energy function, we leverage optimization methods in machine learning to optimize the message update rules, and proposed Momentum-BP and AdaGrad-BP. In addition to improving decoding accuracy, experiments on trapping sets in surface codes show that Momentum-BP and AdaGrad-BP can smooth the message update process and reduce oscillation amplitude.
@@
Furthermore, realizing the importance of a priori and a posteriori probabilities involved in the iteration of BP, we propose EWAInit-BP which dynamically updates the a priori probabilities involved in each iteration by utilizing the a posteriori probabilities output by previous iterations. Experiments on short toric codes, planar surface codes, and XZZX surface codes (under bias noise) show that, under both parallel scheduling and serial scheduling, EWAInit-BP outperforms existing BP improvements without post-processing and outer loops. Notably, the time complexity of all proposed improvements remains $O(j\tau N)$. Thus, EWAInit-BP can be implemented with theoretical constant complexity in parallel while maintaining decoding accuracy.
@@
The structure of this article is as follows. Section II introduces the fundamentals of surface codes and BP decoding, as well as the perspective of BP as a gradient optimization, and summarizes the challenges BP faces in decoding surface codes. Section III introduces Momentum-BP, AdaGrad-BP and EWAInit-BP. The results of the simulation are presented in Section IV. Section V summarizes our work and discusses the future challenges.
II Preliminaries
II.1 Surface codes
@@
A stabilizer code is defined by an abelian subgroup $\mathcal{S}$ of the Pauli group $\mathcal{P}_{n}$ excluding $-I$, composed of operators $S\in\mathcal{P}_{n}$ such that $S|\psi\rangle=|\psi\rangle$ for all codewords $|\psi\rangle$ in the code space. The code space is the simultaneous $+1$ eigenspace of all elements of $\mathcal{S}$. The stabilizers must commute with each other, which means $[S_{i},S_{j}]=0$ for $\forall S_{i},S_{j}\in\mathcal{S}$. The dimension of the code space is determined by the number of independent generators of $\mathcal{S}$. For a stabilizer code with $N$ physical qubits and $K$ logical qubits, $\mathcal{S}$ has $N-K$ independent generators.
@@
The stabilizers are used to detect errors by measuring their eigenvalues. The set of measurement outcomes, known as the error syndrome, contains some information about the errors on physical qubits. The logical operators of a stabilizer code are Pauli operators that commute with all elements of $\mathcal{S}$ but are not in $\mathcal{S}$, thus in $\mathcal{N(S)\backslash S}$ where $\mathcal{N(S)}=\{E:ES=SE,\ \forall S\in\mathcal{S}\}$. They act on the encoded logical qubits without disturbing the code space. Due to this, the logical operators can be also regarded as undetectable errors, in the case that they are not performed on the physical qubits according to our requirements. The code distance $D$ is defined as the minimum weight of logical operators. A quantum stabilizer code can be represented by $[[N,K,D]]$.
@@
The most eye-catching stabilizer codes in recent years might be surface codes. Surface codes are a class of quantum Low-Density Parity-Check (LDPC) codes [39] where qubits are placed on a 2-dimensional lattice, relying on the topology of the surface to protect quantum information [40]. Logical qubits are encoded in global degrees of freedom, and larger lattices provide larger code distances. There are some common properties of surface codes, such as the code distance scales with the code length as $O(\sqrt{N})$, while the code rate is asymptotically zero. In this article, toric codes refer to the ones with closed boundary, and planar surface codes refer to the ones with open boundary. The stabilizers of these two types are all Pauli $X$ or Pauli $Z$. Besides, XZZX surface codes refer to the toric codes whose stabilizers are in the form of $XZZX$.
@@
A $[[2L^{2},2,L]]$ toric code has a 2-dimensional torus topology, and the closed boundaries makes it translationally invariant to the error syndrome, which is beneficial to decoding algorithms (such as BP and neural networks). The Euler characteristic of the torus is 0, but the X-type and Z-type stabilizers of the toric code both have a redundancy, resulting in a total of two encoded logical qubits.
@@
A $[[2L^{2}-2L+1,1,L]]$ planar surface code is defined on a surface with open boundaries, which makes the stabilizers near the boundaries having lower weights. This reduction in stabilizer weight results in a slight decrease in the error correction capability compared to the toric code, however, simplifies its implementation in physical quantum systems. The lattice structure of the planar surface codes are illustrated in Fig. 1.
@@
A $[[L^{2},2,L]]$ XZZX surface code is a non-CSS code designed to handle biased noise more effectively. The stabilizers of these codes have the form of $XZZX$. This alternating pattern of Pauli $X$ and $Z$ operators allows the code to reach the theoretical maximum threshold of 50% under pure Pauli $X$, pure Pauli $Y$ and pure Pauli $Z$ noise.
@@
Figure 1: The lattice of the $[[2L^{2}-2L+1,1,L]]$ planar surface code with L=3. Circles represent data qubits, yellow squares represent X-type stabilizers, green squares represent Z-type stabilizers, and arrows indicate the coupling between stabilizers and qubits.
II.2 Belief propagation decoding
II.2.1 BP over GF(4)
@@
BP algorithm passes messages on a Tanner graph, where stabilizers and qubits are represented as check nodes and variable nodes, respectively. Over GF(4) which this article mainly focuses on, each variable node corresponds to a probability vector of all Pauli errors on one qubit.
Besides, one can use log-likelihood ratio (LLR) as the message passed between variable nodes and check nodes. In this way, belief propagation (abbreviated as LLR-BP4) receives a priori probability vectors describing each type of Pauli error on each qubit $Q^{(0)}=\{p_{v_{j}}^{X},p_{v_{j}}^{Y},p_{v_{j}}^{Z}\}_{j=1}^{N}$, and initializes the variable-to-check messages as [20]
@@
One iteration of BP consists of one horizontal update, one vertical update and one hard decision. In the horizontal update, each check node calculates the feedback based on the messages received from the variable nodes but not from the target variable node in last vertical update, sending a check-to-variable message as follows
@@
In the vertical update, variable nodes aggregate the messages from check nodes and, after a similar extrinsic calculation as above, send variable-to-check messages as follows
@@
Then, for each variable node, if all log-likelihood ratios for Pauli errors are greater than 0, then the hard decision result is I; otherwise, it is the Pauli error with the smallest LLR value. The hard decisions of all qubits combine to form the error estimate $\hat{E}$. If $\hat{E}$ corrects the error syndrome, the algorithm converges; otherwise, it returns to the horizontal update for a new iteration. If the maximum number of iterations is reached, decoding is considered to have failed.
BP can be performed under either parallel or serial scheduling [42]. Under parallel scheduling, all variable nodes and check nodes update their messages simultaneously at each iteration, allowing for hardware parallelization and constant time complexity. On the other hand, the serial scheduling updates the messages of variable nodes and check nodes sequentially. This method leads to faster convergence and slightly higher accuracy, though it has a more complex hardware implementation and higher time complexity.
II.2.2 BP as gradient optimization
@@
From this gradient optimization perspective, MBP [20] adjusts the ”learning rate” of message updates to converge more quickly to the equivalent energy minimum, which is common in degenerate quantum codes.
II.2.3 Hardness of BP in decoding surface codes
The decoding accuracy of BP in decoding surface codes is unsatisfactory due to several reasons, including cycles, degeneracy and trapping sets.
@@ II.2.3 Hardness of BP in decoding surface codes
Short cycles cause the joint probability calculations in BP to be dependent, which violates the independence assumption of updating variable nodes Eq. (7) and check nodes Eq. (5) [19]. Surface codes exhibit 8-cycles over GF(2) and 4-cycles over GF(4), exacerbating this issue.
@@
Optimizing BP from the first two perspectives involves modifying the underlying principles of the algorithm. This article will focus on trapping sets and seek methods to break the symmetry in message passing.
@@
III Belief Propagation Optimization
Surface codes violate the assumptions of BP’s objective function and probability independence, making BP decoding difficult to converge. These issues are manifested in symmetric trapping sets, where the local structure of each node is identical, and message updates are symmetric under specific error patterns. Breaking the former symmetry involves modifying the code structure, while in this section, we focus on breaking the latter, the symmetry of message updates.
III.1 Message update optimization
III.1.1 Update smoothing